SSG Woods deposited $1,500 in his savings account. If the bank pays him 8 1/2 % interest per year, how much money will he make in interest at the end of one year?
Answer by ewatrrr(24785)


Hi,
$1,500 at 8 1/2 % interest per year
8 1/2% = 8.5 % = 8.5/100 = .085 in decimal form
Interest after one year will be = .085*$1500 = $127.50